# Illuminating Black Hole Shadow with Dark Matter Annihilation

### v2

@@ I.2 Dark Matter Spike Annihilation
This section delves into the DM source function $q(\vec{r},p)$, which connects to the DM energy density $\rho(\vec{r}\,)$ via the equation:
@@
In this formula, $m_{\rm DM}$ represents the mass of the DM particles, and $\langle\sigma v\rangle$ denotes their thermally averaged cross-section times the relative velocity. The term ${\rm d}N^{\rm inj}_{e^{\pm},i}/{\rm d}p$ is the electron-positron ($e^{\pm}$) injection spectrum for the annihilation channel indexed by $i$, which has a branching ratio ${\rm BR}_{i}$. For our analysis, we focus on the benchmark annihilation channels $i=b\bar{b}$ and $e^{+}e^{-}$, assuming $100\%$ branching ratios for both. The injection spectra for these channels are sourced from the PPPC Cirelli et al. (2011) and MadDM Ambrogi et al. (2019) packages, with the latter specifically applied to the $e^{+}e^{-}$ channel for DM masses below 55 55 GeV.
The DM energy density profile adopts the spike density model proposed in Ref. Gondolo and Silk (1999) as an illustration. Our investigation specifically concerns the SMBH M87, with a mass of $M_{\rm BH}=6.5\times 10^{9}\,M_{\odot}$ Akiyama et al. (2019a, c), and an associated gravitational radius $r_{g}\equiv GM_{\rm BH}\approx 3.1\times 10^{-4}\,{\rm pc}$. We adopt the parameterization and normalization approach as outlined in Refs. Lacroix et al. (2015, 2017). Beginning with a basic power-law halo profile $\rho_{\rm halo}(r)\propto r^{-1}$, which is representative of the inner regions of an NFW profile Navarro et al. (1996), the resultant spike profile due to adiabatic accretion by the SMBH at the galactic center is described as follows:
@@
Key quantities within this equation are elucidated below:
  • $\gamma_{\rm sp}$ and $r_{\rm sp}$ represent the slope index and the extent of the spike, respectively. These parameters are influenced by the initial halo profile without the SMBH. For an NFW halo profile described by $\rho_{\rm halo}(r)\propto r^{-1}$ Navarro et al. (1996), we adopt $\gamma_{\rm sp}=7/3$ and $r_{\rm sp}=0.001M_{\rm BH}^{3/2}\rho_{0}^{-3/2}r_{0}^{-7/2}$ Gondolo and Silk (1999), where $\rho_{0}$ and $r_{0}$ are scale factors established via the normalization condition. The total DM mass within the SMBH’s sphere of influence can be inferred by jointly analyzing stellar motion Gebhardt et al. (2011) and the size of the photon ring Akiyama et al. (2019a, c), with an associated uncertainty of about $10\%\,M_{\rm BH}$. Accordingly, we adopt the normalization approach from Refs. Lacroix et al. (2015, 2017), setting $r_{0}=10^{5}\,r_{g}$, which is less than $r_{\rm sp}$, and determining the enclosed total DM mass to be $\Delta M_{\rm BH}\approx 5\times 10^{8}M_{\odot}$. The determination of $\rho_{0}$ is guided by the relation:
    $\displaystyle\int\limits_{r\leq r_{0}}{\rm d}^{3}\vec{r}\,\rho(\vec{r}\,)\approx\Delta M_{\rm BH}\,.$ (S18)
    where the integral primarily reflects contributions from larger radii, suggesting that any potential saturation at the inner region negligibly affects the normalization. Consequently, the values of $\rho_{0}$ and $r_{\rm sp}$ are estimated as follows:
    $\displaystyle\rho_{0}$ $\displaystyle\approx\Delta M_{\rm BH}/(6\pi r_{0}^{3})\approx 3.3\times 10^{4}\,{\rm GeV\,cm^{-3}}\left(\frac{\Delta M_{\rm BH}}{5\times 10^{8}\,M_{\odot}}\right)\left(\frac{r_{0}}{10^{5}\,r_{g}}\right)^{-3}\,,$ (S19)
    $\displaystyle r_{\rm sp}$ $\displaystyle\approx 4.0\times 10^{5}\,r_{g}\,\left(\frac{M_{\rm BH}}{6.5\times 10^{9}\,M_{\odot}}\right)^{3/2}\left(\frac{\rho_{0}}{3.3\times 10^{4}\,\mathrm{GeV/cm^{3}}}\right)^{-3/2}\left(\frac{r_{0}}{10^{5}\,r_{g}}\right)^{-7/2}\,.$ (S20)
  • The saturation region, defined by a density $\rho_{\rm sat}\equiv m_{\rm DM}/(\langle\sigma v\rangle\,t_{\rm BH})$, emerges as a consequence of DM annihilations Gondolo and Silk (1999). Here, the age of the BH is considered to be $t_{\rm BH}=10^{8}\,{\rm yr}$ Lacroix et al. (2017). The radius of this saturated region, $r_{\rm sat}=r_{0}\,(\rho_{0}/\rho_{\rm sat})^{1/\gamma_{\rm sp}}$, is calculated to ensure a seamless transition in the density profile. For M87, the calculated values are as follows:
@@
    This relationship implies that a larger cross-section correlates with an expanded saturated region. When $r_{\rm sat}$ falls beneath the critical radius $r_{\rm crit}(\theta,a_{J})$, distortion of DM profile due to annihilation can be considered negligible.
  • $r_{\rm crit}(\theta,a_{J})$ represents the inner boundary for the density profile, uniquely affecting the profile’s spherical symmetry centered around the BH with spin $a_{J}$, due to its dependency on the polar angle $\theta$. For massive particles bound by the BH’s gravitational pull, any orbit dipping below $r_{\rm crit}$ is destined to plunge into the BH. The analytic expression for $r_{\rm crit}$ can be found in Ref. Hod (2013). Specifically, for M87 with a spin parameter $a_{J}=0.9375$, $r_{\rm crit}$ spans from $1.5\,r_{g}$ at the BH’s equatorial plane ($\theta=\pi/2$) to $3.5\,r_{g}$ along the polar axis ($\theta=0$). It is noteworthy that a comprehensive general relativistic analysis would predict a denser spike profile Sadeghian et al. (2013); Ferrer et al. (2017); Zhang and Tang (2024) than the initial model Gondolo and Silk (1999) incorporating the cutoff at $r_{\rm crit}$, underscoring our benchmark selection as a conservative approach for establishing constraints on DM annihilation.
Figure S1: DM spike density profile along the polar axis $(\theta=0)$ outside M87, utilized in this study for different values of the thermally averaged cross-section times velocity, $\langle\sigma v\rangle$, with a fixed DM particle mass of $10\,\mathrm{GeV}$. The spin of the BH that determines the cutoff radius $r_{\rm crit}$ is $a_{J}=0.9375$.
The benchmark DM profile at the polar axis $(\theta=0)$ is shown in Fig. S1.
I.3 Magnetic Field Profiles from GRMHD Simulation
@@
with $(\hat{e}_{x},\hat{e}_{\phi},\hat{e}_{z})$ representing the unit vectors in the corresponding directional coordinates.
To facilitate analysis, we define dimensionless variables:
$\displaystyle\tilde{r}\equiv\sqrt{\tilde{x}^{2}+\tilde{z}^{2}}\equiv\frac{r}{r_{g}},\qquad\tilde{B}\equiv\frac{B}{1\,G}.$ (S24)
The GRMHD normalization ensures that the total synchrotron radiation intensity at 230230 230230 GHz matches 0.60.6 0.60.6 Jy, assuming $R_{\rm high}=1$ and $a_{J}=0.9375$. This condition is detailed further in the subsequent section. Given the expectation for the time-averaged magnetic field to display rotational symmetry, we calculate the averaged magnetic field components over $\phi$ and time using data from GRMHD simulations.
@@
Our analysis divides the magnetic field fitting process into two principal areas: the jet region, characterized by magnetic field lines intersecting the event horizon near the BH’s polar axis, and the disk region, where magnetic field lines are truncated at the BH’s equatorial plane. The fitting techniques for each region are outlined separately below:
@@
  • In the jet region, we adopt a semi-analytic ansatz based on Ref. Anantua et al. (2020) and numerically refit the parameters using GRMHD data. The ansatz implements parabolic functions centered around the polar axis given that the magnetic field configuration in this area is predominantly parabolic. A self-similarity variable, $\xi\equiv\tilde{x}^{2}/\tilde{z}$, is introduced to categorize magnetic field lines via magnetic flux $\Phi(\xi)$ and current $I(\xi)$. By applying Maxwell’s equations within the ideal MHD framework, the magnetic field components can be depicted as:
    $\begin{split}\tilde{B}_{x}&=\tilde{x}\Phi^{\prime}/(2\pi\tilde{z}^{2}),\\ \tilde{B}_{\phi}&=I/(2\pi\tilde{x}),\\ \tilde{B}_{z}&=\Phi^{\prime}/(\pi\tilde{z}),\end{split}$ (S25)
    with $\Phi^{\prime}\equiv{\rm d}\Phi/{\rm d}\xi$. The current satisfies the equation $I=-2|\vec{\Omega}|\xi\Phi^{\prime}$ under force-free conditions and the Ohm’s law, where $\vec{\Omega}$ denotes the angular velocity of the magnetic field line. The ansatz for $\Phi^{\prime}(\xi)$ is given by:
    $\displaystyle\Phi^{\prime}(\xi)=-c_{1}\tanh(c_{2}\cdot\xi)+c_{3},$ (S26)
[…]
    $\displaystyle\left|\vec{\Omega}(\xi,a_{J})\right|$ $\displaystyle=d_{1}+\frac{d_{2}}{\left(\xi+\sqrt{d_{2}/(\Omega_{H}/2-d_{1})}\right)^{2}},$ (S27)
[…]
@@
  • In the disk region, the magnetic field lines within the $x-z$ plane adopt a hyperbolic-like configuration, consequently motivating us to use the following parameterization:
    $\frac{\tilde{B}_{x}}{\tilde{B}_{z}}=\frac{\tilde{z}\cdot K(\tilde{r})}{\tilde{x}},$ (S28)
    where $K(\tilde{r})$ represents the distortion near the event horizon and is modeled through a polynomial function: $K(\tilde{r})=1000\,\tilde{r}^{-4}+100\,\tilde{r}^{-3}+8\,\tilde{r}^{-2}+4\,\tilde{r}^{-1}+1$. This equation features two groups of hyperbolas centered on the $x$-axis and $z$-axis, respectively; however, for our analysis focusing on the disk region, only the hyperbolas focused on the $x$-axis are considered. The separation between the two groups is identified as the boundary between the jet and the disk.
    The overall magnitude of the magnetic field and its $\phi$-component are determined by:
    $\begin{split}\tilde{B}&=\left[b_{1}-b_{2}\exp(b_{3}(\theta-\pi/4))\right]\tilde{r}^{-b_{4}},\\ \tilde{B}_{\phi}&=\left[\frac{a_{1}}{\tilde{r}+a_{2}/(\cos(\theta-\pi/4)-a_{3})}+a_{4}\right]\tilde{B},\end{split}$ (S29)
    where $\theta=\arctan(\tilde{x}/\tilde{z})$. The coefficients determined for the best fit are:
[…]
    As indicated, the $\phi$ component’s proportion, $\tilde{B}_{\phi}/\tilde{B}$, diminishes with increasing $\tilde{r}$. The inclusion of a constant term, $a_{4}$, ensures that this ratio retains a specific value. The magnetic field lines exhibit a spiral pattern around the polar axis in the inner region. The selected coefficients guarantee that the $\tilde{B}_{\phi}$ fraction remains below unity within $2.5\,r_{g}$. The field strength shows a rapid decay with distance from the BH. The coefficient $b_{1}$ parameterizes the strong magnetic field magnitude near the BH.
@@ I.4 Bulk Velocity of Electron-Positron Flows
Analogous to the magnetic field profile delineated in the previous subsection, the bulk velocity $\vec{v}_{b}$ exhibits distinct characteristics in the jet and disk regions, primarily due to the rotation of magnetic field lines within the jet.
Within the disk region, the gravitational influence of the SMBH directs electrons and positrons towards the BH, guiding them along the magnetic field lines. Since the Lorentz force does not affect the kinetic energy of a charged particle, the gravitational pull predominantly determines the bulk velocity’s magnitude. This is represented by the radial infall velocity $v_{b}=(2\,r_{g}/r)^{1/2}$. Consequently, the direction of the bulk velocity aligns with that of the magnetic field line, with its radial component consistently oriented towards the BH.
In the jet region, the magnetic field lines stretch to the event horizon, influenced by the BH rotation, which imparts a drag effect. This interaction induces rotational motion around the BH’s polar axis at an angular velocity $\vec{\Omega}$. Within the frame co-rotating with the magnetic field lines, the electron-positron flow dynamics are shaped by both the inertial centrifugal potential and the BH’s gravitational pull Pu et al. (2017). Assuming a consistent angular velocity $\vec{\Omega}$ for each magnetic field line in the jet, as derived from Eq. (S27), the bulk velocity $\vec{v}_{b}$ is ascertained via energy conservation:
$\begin{split}\frac{{\rm d}\gamma_{v_{b}}}{{\rm d}s}&=\hat{v}_{b}(\vec{r}\,)\cdot\left(-\frac{r_{g}\gamma_{v_{b}}}{r^{3}}\vec{r}-\gamma_{v_{b}}\vec{\Omega}\times\left(\vec{\Omega}\times\vec{r}\right)\right)\\ &=\gamma_{v_{b}}\left(\frac{r_{g}}{r^{2}}\frac{1+\cos^{2}\theta}{\sqrt{1+3\cos^{2}\theta}}-\Omega^{2}r\sin^{2}\theta\right)\,,\end{split}$ (S31)
where $\gamma_{v_{b}}=1/\sqrt{1-v_{b}^{2}}$ signifies the Lorentz factor of the bulk velocity. The stagnation surface, defined where $\mathrm{d}\gamma_{v_{b}}/\mathrm{d}s=0$, is identified by:
$\theta_{\rm stag}(r)=\arccos\left(\sqrt{\frac{\Omega^{2}r^{3}-r_{g}}{\Omega^{2}r^{3}+r_{g}}}\right),$ (S32)
dividing the space into regions where the bulk velocity either accelerates towards the BH (inner region) or outward towards infinity (outer region). At this surface, the bulk velocity is set to zero, $\vec{v}_{b}=0$. By defining an effective potential:
@@
Eq. (S31) simplifies to ${\rm d}\ln\gamma_{v_{b}}=-{\rm d}V_{\rm eff}(r,\theta)$. Consequently, the bulk velocity magnitude inside the stagnation surface is calculated as:
@@
where $(r_{\rm stag},\theta_{\rm stag})$ denotes the intersection of the magnetic field line, traversing through $(r,\theta)$, with the stagnation surface.

### v3

@@ I.2 Dark Matter Spike Annihilation
This section examines the DM annihilation source function $q(\vec{r},p)$, which is proportional to the square of the DM energy density $\rho(\vec{r}\,)$:
@@
Here, $m_{\rm DM}$ is the DM particle mass, and $\langle\sigma v\rangle$ denotes the thermally averaged annihilation cross-section times the relative velocity. ${\rm d}N^{\rm inj}_{e^{\pm},i}/{\rm d}p$ represents the electron-positron ($e^{\pm}$) injection spectrum for the annihilation channel indexed by $i$, weighted by its branching ratio ${\rm BR}_{i}$. We focus on two benchmark annihilation channels, $i=b\bar{b}$ and $e^{+}e^{-}$, assuming a $100\%$ branching ratio for each. The injection spectra for these channels are obtained from the PPPC Cirelli et al. (2011) and MadDM Ambrogi et al. (2019) packages, with the latter specifically applied to the $e^{+}e^{-}$ channel for DM masses below $5$ GeV.
For the DM energy density profile, we adopt the spike density model proposed in Ref. Gondolo and Silk (1999) as an illustrative example. Our analysis specifically focuses on the SMBH M87, with a mass of $M_{\rm BH}=6.5\times 10^{9}\,M_{\odot}$ Akiyama et al. (2019a, d) and a gravitational radius of $r_{g}\equiv GM_{\rm BH}\approx 3.1\times 10^{-4}\,{\rm pc}$. We follow the parameterization and normalization approach outlined in Refs. Lacroix et al. (2015, 2017).
Starting with a basic power-law halo profile, $\rho_{\rm halo}(r)\propto r^{-1}$, which approximates the inner regions of a Navarro-Frenk-White (NFW) profile Navarro et al. (1996), the resulting spike profile—formed due to adiabatic accretion by the SMBH at the galactic center—is given by:
@@
The parameters in this equation are explained below:
  • The parameters $\gamma_{\rm sp}=7/3$ and $r_{\rm sp}=0.001M_{\rm BH}^{3/2}\rho_{0}^{-3/2}r_{0}^{-7/2}$ represent the slope index and the extent of the DM spike, respectively, which are determined by the initial halo profile in the absence of the SMBH Gondolo and Silk (1999). Here, we assume an NFW halo profile, given by $\rho_{\rm halo}(r)\propto r^{-1}$ Navarro et al. (1996). The normalization factors are chosen as $\rho_{0}=3.3\times 10^{4}\,{\rm GeV\,cm^{-3}}$ and $r_{0}=10^{5}\,r_{g}$ to ensure that the total enclosed DM mass within $r_{\rm sp}$ is $\Delta M_{\rm BH}\approx 5\times 10^{8}M_{\odot}$. This normalization approach follows Refs. Lacroix et al. (2015, 2017), incorporating approximately $10\%$ uncertainty in the BH mass measurement, which arises from joint analyses of stellar motion Gebhardt et al. (2011) and the size of the photon ring Akiyama et al. (2019a, d).
  • The saturation region, characterized by a density of $\rho_{\rm sat}\equiv m_{\rm DM}/(\langle\sigma v\rangle\,t_{\rm BH})$, arises due to DM annihilation Gondolo and Silk (1999). Here, we assume a BH age of $t_{\rm BH}=10^{8}\,{\rm yr}$ Lacroix et al. (2017). The radius of this region, given by $r_{\rm sat}=r_{0}\,(\rho_{0}/\rho_{\rm sat})^{1/\gamma_{\rm sp}}$, is determined to ensure a smooth transition in the density profile. For M87, the values are:
@@
    This relation indicates that a larger annihilation cross-section leads to a more extended saturation region. If $r_{\rm sat}$ is smaller than the critical radius $r_{\rm crit}(\theta,a_{J})$, the impact of annihilation on the DM profile becomes negligible.
  • The parameter $r_{\rm crit}(\theta,a_{J})$ defines the inner cutoff boundary of the DM density profile. For massive particles bound by the BH gravitational potential, any orbit that dips below $r_{\rm crit}$ inevitably plunges into the BH. The dimensionless quantity $\mathcal{R}_{\rm crit}\equiv r_{\rm crit}/r_{g}$ is determined as the largest real root of the equation Hod (2013):
    $\mathcal{R}_{\rm crit}^{4}-4\mathcal{R}_{\rm crit}^{3}-a_{J}^{2}(1-3\cos^{2}\theta)\mathcal{R}_{\rm crit}^{2}+a_{J}^{4}\cos^{2}\theta+4a_{J}\sqrt{(1-\cos^{2}\theta)(\mathcal{R}_{\rm crit}^{5}-a_{J}^{2}\mathcal{R}_{\rm crit}^{3}\cos^{2}\theta)}=0$ (S20)
    which is non-spherical and depends on both the polar angle $\theta$ and the BH spin parameter $a_{J}$.
    For a BH with $a_{J}=0.9375$ ($0.5$), the critical radius $r_{\rm crit}$ ranges from $1.5$ ($2.9$) $r_{g}$ at the equatorial plane ($\theta=\pi/2$) to $3.5$ ($3.9$) $r_{g}$ along the polar axis $\theta=0$. A full general relativistic analysis predicts a denser spike profile Sadeghian et al. (2013); Ferrer et al. (2017); Zhang and Tang (2024) than the initial model Gondolo and Silk (1999) near the cutoff radius $r_{\rm crit}$, reinforcing our benchmark choice as a conservative approach for setting constraints on DM annihilation.
Figure S1: DM density profile outside the SMBH M87, with a gravitational radius of $r_{g}=3.1\times 10^{-4}$ pc, as adopted in this study. The profile follows the DM spike model Gondolo and Silk (1999), transitioning from saturation at $\rho_{\rm sat}$ to an outer power-law behavior described by the NFW profile Navarro et al. (1996), ensuring a smooth density distribution. The profiles are shown for different values of the thermally averaged annihilation cross-section $\langle\sigma v\rangle$, while fixing the DM particle mass at $10\,\mathrm{GeV}$. A larger annihilation cross-section or lower DM mass results in a lower saturation density in the inner region. The inner cutoff radius, $r_{\rm crit}$, is determined by the BH spin parameter and polar angle, set here to $a_{J}=0.9375$ and $\theta=0$, yielding $r_{\rm crit}\approx 3.5\,r_{g}$.
The benchmark DM profile outside the SMBH M87 ($r_{g}=3.1\times 10^{-4}$ pc) is shown in Fig. S1, where we adopt the cutoff radius $r_{\rm crit}$ at the polar axis $(\theta=0)$ for a BH spin of $a_{J}=0.9375$.
I.3 Magnetic Field Profiles from GRMHD Simulations
@@
with $(\hat{e}_{x},\hat{e}_{\phi},\hat{e}_{z})$ representing the unit vectors in the corresponding directional coordinates. To facilitate analysis, we define dimensionless variables:
$\displaystyle\mathcal{R}\equiv\sqrt{\mathcal{X}^{2}+\mathcal{Z}^{2}}\equiv\frac{r}{r_{g}},\qquad\mathcal{B}\equiv\frac{B}{1\,G}.$ (S22)
In our analysis, we consider both $a_{J}=0.9375$ and $a_{J}=0.5$ to account for uncertainties in the BH spin (see Sec. IV.1). The GRMHD profiles are normalized such that the total synchrotron radiation intensity at $230$ GHz matches the observed value of $0.6$ Jy. Assuming that the time-averaged magnetic field exhibits approximate rotational symmetry, we compute the azimuthal and temporal averages of the magnetic field components using GRMHD simulation data.
@@
We divide the magnetic field fitting procedure into two main regions: the jet region, defined by magnetic field lines that thread the event horizon near the polar axis, and the disk region, where the field lines terminate at the equatorial plane of the BH. The fitting methodology for each region is described separately below.
@@
  • In the jet region, we adopt a general analytic jet model based on Refs. Pu et al. (2017); Pu and Takahashi (2020) and numerically refit the magnetic field strengths using GRMHD simulation data. The magnetic field configuration in this region is predominantly parabolic, and the model employs parabolic streamlines centered around the polar axis. A self-similarity variable $\Psi\equiv\mathcal{R}(1-\cos\theta)$ is introduced to label magnetic field lines through the magnetic flux $\Phi(\Psi)$ and the field line angular velocity $\Omega_{F}(\Psi)$.
    By applying Maxwell’s equations within the ideal MHD framework, the magnetic field components can be depicted as:
    $\begin{split}&\mathcal{B}_{x}=\mathcal{X}\Phi^{\prime}/[2\pi\mathcal{R}(\mathcal{Z}+\mathcal{R})],\\ &\mathcal{B}_{\phi}=\Omega_{F}g_{\phi\phi}\Phi^{\prime}/(2\pi\mathcal{R}),\\ &\mathcal{B}_{z}=\Phi^{\prime}/(2\pi\mathcal{R}),\end{split}$ (S23)
    where $\Phi^{\prime}\equiv{\rm d}\Phi/{\rm d}\Psi$, and $g_{\phi\phi}$ is a component of the Kerr metric of the BH.
    The angular velocity of the magnetic field lines, $\Omega_{F}$, is given by Pu and Takahashi (2020):
    $\displaystyle\Omega_{F}(\Psi,a_{J})=\frac{\sin^{2}\theta_{H}(1+\ln\mathcal{G})}{4\,\ln 2+\sin^{2}\theta_{H}+(\sin^{2}\theta_{H}-2\mathcal{G})\ln\mathcal{G}}\Omega_{H}.$ (S24)
    Here, $\Omega_{H}=a_{J}/(2\mathcal{R}_{H})$ denotes the angular velocity of the BH outer horizon and $\mathcal{R}_{H}=1+(1-a_{J}^{2})^{1/2}$. The factor $\mathcal{G}=1+\cos\theta_{H}$ depends on the polar angle $\theta_{H}$ of the magnetic field at the horizon, defined via $\theta_{H}(\Psi)\equiv\arccos(1-\Psi/\mathcal{R}_{H})$. The jet region corresponds to the range $0\leq\Psi<\mathcal{R}_{H}$, with its boundary given by $\mathcal{R}_{H}=\mathcal{R}(1-\cos\theta)$.
    To match the magnetic field strength from GRMHD simulations, we use an ansatz for $\Phi^{\prime}(\Psi)$:
    $\displaystyle\Phi^{\prime}(\Psi)=c_{1}\exp(-c_{2}\cdot\Psi),$ (S25)
    yielding the best-fit coefficients:
    $\left(c_{1},\,c_{2}\right)=\begin{cases}\left(123,\,0.679\right),\quad&a_{J}=0.9375,\\ \left(126,\,0.345\right),\quad&a_{J}=0.5.\end{cases}$ (S26)
@@
  • In the disk region, the magnetic field lines within the $x$$z$ plane exhibit a hyperbolic-like structure. This motivates the following parameterization:
    $\frac{\mathcal{B}_{x}}{\mathcal{B}_{z}}=\frac{\mathcal{Z}\cdot K(\mathcal{R})}{\mathcal{X}},$ (S27)
    where $K(\mathcal{R})$ accounts for distortions near the event horizon and is modeled as a polynomial: $K(\mathcal{R})=1000\,\mathcal{R}^{-4}+100\,\mathcal{R}^{-3}+8\,\mathcal{R}^{-2}+4\,\mathcal{R}^{-1}+1$. This formulation describes two branches of hyperbolas, one centered on the $x$-axis and the other on the $z$-axis. In our analysis, which focuses on the disk region, we retain only the branch centered on the $x$-axis. The separation between the two branches defines the boundary of the disk region, determined by solving Eq. (S27) under the constraint that the hyperbolic solution intersects the origin of the $x$$z$ plane.
    The overall magnetic field strength and its azimuthal component are fit using:
    $\begin{split}\mathcal{B}_{2D}&\equiv\sqrt{\mathcal{B}_{x}^{2}+\mathcal{B}_{z}^{2}}=b_{1}\mathcal{R}^{b_{2}\sin(4\theta-b_{3})-b_{4}},\\ \mathcal{B}_{\phi}&=a_{1}\mathcal{R}^{a_{2}\sin(4\theta-a_{3})-a_{4}}.\end{split}$ (S28)
    The best-fit parameters for the two spin cases are:
    $\left(a_{1},\,a_{2},\,a_{3},\,a_{4}\right)=\begin{cases}\left(38.1,\,0.511,\,2.34,\,1.81\right),\quad&a_{J}=0.9375,\\ \left(19.4,\,0.395,\,2.72,\,1.56\right),\quad&a_{J}=0.5.\end{cases}$ (S29)
    $\left(b_{1},\,b_{2},\,b_{3},\,b_{4}\right)=\begin{cases}\left(52.1,\,0.579,\,2.39,\,2.29\right),\quad&a_{J}=0.9375,\\ \left(94.1,\,0.596,\,2.70,\,2.39\right),\quad&a_{J}=0.5.\end{cases}$ (S30)
    As expected, the fraction $\mathcal{B}_{\phi}/\mathcal{B}$ increases with BH spin.
@@ I.4 Bulk Velocity of Electron-Positron Flows
Analogous to the magnetic field configuration discussed in the previous subsection, the bulk velocity $\vec{v}_{b}$ exhibits distinct behaviors in the jet and disk regions, primarily due to the rotation of magnetic field lines in the jet.
In the disk region, the SMBH’s gravitational pull drives electrons and positrons inward along magnetic field lines. As a result, the direction of the bulk velocity aligns with the field lines, with a radial component consistently directed toward the BH. Since the Lorentz force does not change the kinetic energy of a charged particle, the magnitude of the bulk velocity is predominantly governed by gravity and approximated by the radial infall velocity $v_{b}=(2\,r_{g}/r)^{1/2}$.
In the jet region, magnetic field lines extend to the event horizon and co-rotate with the BH due to frame dragging. This rotation imparts angular momentum to the plasma, inducing rotational motion around the BH’s polar axis with angular velocity $\vec{\Omega}$. In the co-rotating frame, the electron-positron dynamics are influenced by both the BH’s gravity and the inertial centrifugal potential Pu et al. (2017); Pu and Takahashi (2020). Assuming a constant angular velocity $\vec{\Omega}\equiv\Omega_{F}(\Psi)\hat{e}_{z}$ along each magnetic field line, as given in Eq. (S24), the bulk velocity $\vec{v}_{b}$ is determined from energy conservation:
$\frac{{\rm d}\gamma_{v_{b}}}{{\rm d}s}=\hat{v}_{b}(\vec{r}\,)\cdot\left(-\frac{r_{g}\gamma_{v_{b}}}{r^{3}}\vec{r}-\gamma_{v_{b}}\vec{\Omega}\times\left(\vec{\Omega}\times\vec{r}\right)\right)\,$, (S31)
where $\gamma_{v_{b}}=(1-v_{b}^{2})^{-1/2}$ is the Lorentz factor associated with the bulk motion.
The stagnation surface—where $\mathrm{d}\gamma_{v_{b}}/\mathrm{d}s=0$—marks the transition between inflow and outflow, and is given by:
$r_{\rm stag}(\Psi)=\frac{r_{g}}{\Omega_{F}(\Psi)\sqrt{\Psi}}\,$, (S32)
such that the bulk velocity vanishes, $\vec{v}_{b}=0$, at this location.
By defining an effective potential:
@@
Eq. (S31) simplifies to: ${\rm d}\ln\gamma_{v_{b}}=-{\rm d}V_{\rm eff}(r,\theta)$. Integrating this relation yields the bulk velocity magnitude on either side of the stagnation surface:
@@
where $(r_{\rm stag},\theta_{\rm stag})$ corresponds to the point at which the magnetic field line passing through $(r,\theta)$ intersects the stagnation surface.